OpenStudy (anonymous):

What is the equation of a line which includes points (-5, 5) and (2, 5)?

OpenStudy (igreen):

First find the slope. \(\sf m = \dfrac{y_2-y_1}{x_2-x_1}\) \(\sf m = \dfrac{5-5}{2-(-5)}\) Simplify

OpenStudy (anonymous):

7?

OpenStudy (igreen):

Yes.. 2 - (-5) = 7..what about 5 - 5?

OpenStudy (anonymous):

0

OpenStudy (igreen):

Yes, now divide 0 / 7